# Multi-Agent Reinforcement Learning for Task Offloading in Wireless Edge Networks

### v1

@@ Abstract
In edge computing systems, autonomous agents must make fast local decisions while competing for shared resources. Existing MARL methods often resume to centralized critics or frequent communication, which fail under limited observability and communication constraints. We propose a decentralized framework in which each agent solves a constrained Markov decision process (CMDP), coordinating implicitly through a shared constraint vector. For the specific case of offloading, e.g., constraints prevent overloading shared server resources. Coordination constraints are updated infrequently and act as a lightweight coordination mechanism. They enable agents to align with global resource usage objectives but require little direct communication. Using safe reinforcement learning, agents learn policies that meet both local and global goals. We establish theoretical guarantees under mild assumptions and validate our approach experimentally, showing improved performance over centralized and independent baselines, especially in large-scale settings.
@@ 4 Using MARL for optimizing offloading
We remark that DCC is not a single algorithm but a general framework compatible with a broad class of reinforcement learning methods for local MDPs, including value-based and policy-gradient approaches. In this section we outline its formal structure and learning process across three timescales, with full details reported in Appendix A.
@@
Together, these components yield a principled and practical algorithm for decentralized coordination under shared resource constraints, with theoretical support for its approximation quality and optimization dynamics. The technical description of the algorithm is deferred to appendix A.
5 Numerical experiments
Scalability properties
Offloading action frequency
Additional experimental results are reported in the appendix. These include the effect of constraint initialization, comparisons with alternative baselines, the impact of non-linear penalties, and evaluations of some of the theoretical results using linear programming. Together, these supplementary experiments further support the conclusions drawn from the main results.
6 Conclusions
Appendix A DCC framework
This section provides a more precise and formal definition of the DCC framework, elaborating on the key components and methodologies involved in its design and implementation. It is organized into the following steps:
  • Introduction of a decomposable approximation of the immediate reward (Section A.1): we propose a decomposable approximation of the immediate reward that approximates the long-term reward with provably bounded approximation error—and exact equivalence when the congestion function is linear.
  • CMDP of the system (Section A.2): we introduce the single agent CMDP that will by used by the three timescale algorithm
  • Computing the optimal constrained policy (Section A.3): We describe the process of computing the optimal constrained policy for a single agent, given a fixed value of the constraint.
  • Differentiability of the objective (Section A.4): We establish that the long term reward of each agent is differentiable with respect to $\theta$, which allows us to efficiently identify the minimum in (4)
  • Efficient computation of the gradient (Section A.5): We show how the structure of the gradient can be leveraged to accelerate the computation of the infimum of the objective function.
  • some technical proofs conclude the section (sections A.6 and A.7).
A.1 Approximating the objective function via decomposition
@@
meaning that each agent can act independently while still achieving the globally optimal outcome [19, 24]. However, in the setting studied in this work, this condition does not hold due to a coupling term that depends on how many agents select the crowded action. Specifically, the global reward takes the form
@@
$\hat{r}(s,a;\theta)$ $=\sum_{i}f_{i}(s_{i})+\mathbb{I}_{a=a_{\text{crowd}}}(a_{i})\cdot d\left(1+\sum_{j\neq i}\theta_{j}\right)$ (6)
@@
The following result gives an indication of the error that one could have when considering the approximated definition of the reward in (6). In particular, we define $J_{\pi}(\beta)$ the discounted reward associated to a policy $\pi$ when considering the reward $r$ and $\tilde{J}_{\pi}(\beta)$ the one obtained when considering the approximated reward $\hat{r}$ defined in (6).
@@
$\mid J_{\pi}(\beta)-\tilde{J}_{\pi}(\beta)\mid\leq\frac{1}{1-\gamma}\sum_{i}\theta_{i}\left(\frac{\theta_{-i}}{N_{agents}-1}d(N_{agents})-d(1+\theta_{-i})\right)$
@@
$J_{\pi_{\text{composed}}}(\beta)=\hat{J}_{\pi_{\text{composed}}}(\beta)=\sum_{i}\tilde{J}_{i}(\theta^{*},\beta)=J(\beta).$
A.2 Constrained MDP for Each Agent
As introduced in eq. 4, we approximate the global objective by decomposing it into a collection of independent Constrained Markov Decision Processes (CMDPs), one per agent. Constrained Markov Decision Process (CMDP) [3] extends the standard MDP by incorporating constraints on long-term cost signals. In this section, we describe the structure of the CMDP associated with each agent $i$.
@@ A.2 Constrained MDP for Each Agent
The agent receives a reward according to the approximated formulation described in section A.1, where the influence of other agents is captured through the fixed (in the slow and intermediate timescale) parameter $\theta_{-i}$:
$\hat{r}_{i}(s,a;\theta_{-i})=f_{i}(s_{i})+\mathbb{I}_{a_{i}=a_{\text{crowd}}}\cdot d\left(1+\theta_{-i}\right).$
@@
A.3 Learning the Optimal Local Constrained Policy
@@
In particular, we follow the structure of RCPO [29] for constrained optimization, employing either PPO [25], DQN [21] or Q-learning [30], depending on the scale of the system, as the algorithm to handle the fastest timescale
@@
$r_{i}^{\text{total}}(s_{i},a_{i};\theta_{-i})=\hat{r}_{i}(s_{i},a_{i};\theta_{-i})+\lambda_{i}c_{i}(s_{i},a_{i}),$ (10)
where $\hat{r_{i}}$ and $c_{i}$ have been defined in section A.2.
@@
A.4 Coordination via optimization of the constraints
@@
In particular, we want to show that we can converge to the local optimal value of the constraint vector $\theta\in\mathbb{R^{N}}$. The safe RL algorithms described in section A.3 will then be able to find the optimal policy for that value.
We consider the long term Lagrangian reward of each agent, $\tilde{J}^{\ell}_{i}(\theta_{i},\lambda_{i},\pi_{i})$, defined as
$\tilde{J}_{i}^{\ell}(\theta,\lambda_{i},\pi_{i})=\mathbb{E}_{a\sim\pi_{i}(s)}\bigg{[}\sum_{t=0}^{\infty}\gamma^{t}\sum_{t=0}^{T}\hat{r}_{i}(s_{t},a_{t};\theta_{-i})+\lambda(c_{i}(s_{t},a_{t})-\theta_{i})\bigg{]}$ (12)
and the function we want to maximize is
$\tilde{J}^{\ell}(\theta,\lambda,\pi)=\sum_{i}\tilde{J}^{\ell}_{i}(\theta,\lambda_{i},\pi_{i})$ (13)
This term is equivalent to the Lagrangian function in (11), with the difference that the constraint vector $\theta$ is now a variable of the function and not a fixed parameter. From now on we will drop $\pi_{i}$ and $\lambda_{i}$ from the notation for clarity’s sake.
@@
$\tilde{J}^{\ell}_{i}(\theta)$ is differentiable in $\theta$ almost everywhere.
@@
The proof of this result is provided in section A.7, where we show that a variation of the Envelope Theorem [20] can be applied within the context of our setting. ∎
The following result formalizes the intuition that relaxing the constraint for one device decreases its reward (i.e. it gets better from the selfish point of view) and penalizes the others. We also numerically verified these results in section C.7, using linear programming to compute the optimal policy in small systems.
Proposition 2.
The following is verified:
  1. 1.
    $\frac{\partial\tilde{J}_{i}(\theta)}{\partial\theta_{i}}=-\frac{\lambda^{\star}}{M}\leq 0$ a.e., where $\lambda^{\star}$ is the optimal value of the Lagrange multiplier in $\theta_{i}$ and $M$ is a positive constant,
  2. 2.
    $\frac{\partial\tilde{J}_{i}(\theta)}{\partial\theta_{j}}=\theta_{i}\frac{\partial}{\partial\theta_{j}}d(1+\theta_{-i})\geq 0,\ \forall j\neq i$ a.e., where $\pi^{\star}_{i}$ is the optimal policy for agent $i$.
Proof.
In order to discuss the optimization of the virtual constraint we want to use the stochastic gradient ascent methods of the Kiefer-Wolfowitz family [13]. The iteration scheme is
@@
$\left(\hat{g}_{n}\right)_{i}=\frac{\tilde{J}^{\ell}\left(\theta+c_{n}\left(\Delta_{n}\right)_{i}\right)-\tilde{J}^{\ell}\left(\theta\right)}{c_{n}(\Delta_{n})_{i}}$ (15)
@@
Then the algorithm converges to a local optimal value of $\min_{\theta}\sum_{i}\tilde{J}_{i}(\theta,\beta)$.
@@
Lemma 2 guarantees the differentiability of the objective function. The assumptions on the sequences $\alpha_{n}$ and $c_{n}$ allow us to use Proposition 1 in [14] for a subset of $\Theta$ which contains the initial vector $\theta_{0}$. This proves the convergence of the algorithm to a locally optimal solution. ∎
See Proposition 1 in [14] for the statement and the necessary conditions for the convergence to the global optimal point.
A.5 Numerical optimization of the constraints via evaluations of the local policies
@@
$\frac{\partial}{\partial\theta_{i}}\tilde{J}^{\ell}(\theta_{i},\theta_{-i})=\frac{\partial\tilde{J}^{\ell}_{i}(\theta_{i},\theta_{-i})}{\partial\theta_{i}}+\sum_{j\neq i}\frac{\partial\ \tilde{J}^{\ell}_{j}(\theta_{i},\theta_{-i})}{\partial\theta_{-j}}$ (16)
@@
Equation 16 illustrates how the problem structure, together with the form of the constrained value function, can be leveraged to compute the gradient of the approximate objective function efficiently. Specifically, this result allows each component of the gradient at a given point $\theta$ to be estimated using only three stochastic evaluations per agent, rather than $N+1$.
  • $\tilde{J}_{i}(\hat{\theta}_{i},\hat{\theta}_{-i})$
  • $\tilde{J}^{\ell}_{i}(\hat{\theta}_{i}+\epsilon,\hat{\theta}_{-i})$
  • $\tilde{J}^{\ell}_{i}(\hat{\theta}_{i},\hat{\theta}_{-i}+\epsilon)$
In this context, computing $\tilde{J}^{\ell}_{i}(\hat{\theta}_{i},\hat{\theta}_{-i})$ involves finding the optimal constrained policy for agent $i$ under the constraint vector $\hat{\theta}$ using the safe reinforcement learning algorithm described in section A.3, and then evaluating the resulting policy.
Building on Proposition 2, the derivative of the objective could be computed directly, avoiding additional evaluations of the optimal noisy policy and thus improving efficiency. We leave this refinement to future work.
@@
A.6 Proof of Lemma 1
@@
$R_{\pi}(\beta)-\hat{R}_{\pi}(\beta)$ $=\mathbb{E}_{a_{t}\sim\pi}\left[\sum_{t}\gamma^{t}\sum_{i}r_{i}(s_{t},a_{t})\mid s_{0}\sim\beta\right]-\mathbb{E}_{a_{t}\sim\pi}\left[\sum_{t}\gamma^{t}\sum_{i}\hat{r}_{i}(s_{t},a_{t})\mid s_{0}\sim\beta\right]$
@@
$R_{\pi}(\beta)-\hat{R}_{\pi}(\beta)\leq\frac{1}{1-\gamma}\sum_{i}\theta_{i}\left(\frac{\theta_{-i}}{N_{agents}-1}d(N_{agents})-d(1+\theta_{-i})\right)$
@@
$\left|R_{\pi}(\beta)-\hat{R}_{\pi}(\beta)\right|\leq\frac{1}{1-\gamma}\sum_{i}\theta_{i}\left(\frac{\theta_{-i}}{N_{agents}-1}d(N_{agents})-d(1+\theta_{-i})\right)$ (18)
A.7 Proof of lemma 2
Theoretical background

### v2

@@ Abstract
In edge computing systems, autonomous agents must make fast local decisions while competing for shared resources [^1] The code used in our experiments can be found at https://github.com/Andrea-Fox/multiAgentTaskOffloading. [^2] This work has been partially supported by the French National Research Agency (ANR) within the PARFAIT project (ANR-21-CE25-0013).. Existing MARL methods often resume to centralized critics or frequent communication, which fail under limited observability and communication constraints. We propose a decentralized framework in which each agent solves a constrained Markov decision process (CMDP), coordinating implicitly through a shared constraint vector. For the specific case of offloading, e.g., constraints prevent overloading shared server resources. Coordination constraints are updated infrequently and act as a lightweight coordination mechanism. They enable agents to align with global resource usage objectives but require little direct communication. Using safe reinforcement learning, agents learn policies that meet both local and global goals. We establish theoretical guarantees under mild assumptions and validate our approach experimentally, showing improved performance over centralized and independent baselines, especially in large-scale settings.
@@ 4 Using MARL for optimizing offloading
We remark that DCC is not a single algorithm but a general framework compatible with a broad class of reinforcement learning methods for local MDPs, including value-based and policy-gradient approaches. We begin by providing a high-level intuition of the proposed algorithm. A detailed description of each component is then developed in sections 4.1, 4.2, 4.3, 4.4 and 4.5, including the underlying reward approximation, the CMDP formulation, and the multi-timescale optimization process.
@@
Together, these components yield a principled and practical algorithm for decentralized coordination under shared resource constraints, with theoretical support for its approximation quality and optimization dynamics.
Algorithm 1 DCC Framework (Simplified Overview)
1:Initialize: constraint vector $\theta^{0}$, multipliers $\lambda_{i}^{0}$, policies $\pi_{i}^{0}$ for each agent $i$
2:function OptimizeLocalCMDP($i,\theta,n$)
3:  for $m=0,1,\dots,M(n)$ do $\triangleright$ Intermediate timescale: Lagrange multiplier update
4:   for $t=0,1,\dots,T(m,n)$ do $\triangleright$ Fast timescale: policy optimization
5:     Update policy $\pi_{i}$ via RL step on shaped reward $r_{i}^{\text{tot}}$
6:   end for
7:   Update multiplier $\lambda_{i}$
8:  end for
9:  $\hat{J}_{i}(\theta)\leftarrow$ evaluate final policy $\pi_{i}^{\star}$
10:  return $\hat{J}_{i}(\theta)$, final policy $\pi_{i}^{\star}$, multiplier $\lambda_{i}^{\star}$
11:end function
12:
13:
14:for $n=0,1,2,\dots$ do $\triangleright$ Slowest timescale: constraint optimization
15:  for each agent $i=1,\dots,N$ do
16:   $\hat{J}_{i}(\theta),\pi_{i}^{\star},\lambda_{i}^{\star}\leftarrow$ OptimizeLocalCMDP($i,\theta^{n},n$)
17:   Estimate local gradient $\nabla_{\theta}\hat{J}_{i}(\theta^{n})$
18:  end for
19:  Update shared constraint vector: $\theta^{n+1}\leftarrow\theta^{n}+\eta_{n}\nabla\hat{J}(\theta^{n})$
20:end for
A simplified pseudocode for DCC is found in algorithm 1, while the full pseudocode in in appendix D. In the remainder of the section, we investigate more deeply the DCC framework through the following steps:
  • Introduction of a decomposable approximation of the immediate reward (section 4.1): we propose a decomposable approximation of the immediate reward that approximates the long-term reward with provably bounded approximation error—and exact equivalence when the congestion function is linear.
  • CMDP of the system (section 4.2): we introduce the single agent CMDP that will by used by the three timescale algorithm
  • Computing the optimal constrained policy (section 4.3): We describe the process of computing the optimal constrained policy for a single agent, given a fixed value of the constraint.
  • Differentiability of the objective (section 4.4): We establish that the long term reward of each agent is differentiable with respect to $\theta$, which allows us to efficiently identify the minimum in (4)
  • Efficient computation of the gradient (section 4.5): We show how the structure of the gradient can be leveraged to significantly accelerate the computation of the infimum of the objective function.
4.1 Approximating the objective function via decomposition
@@
meaning that each agent can act independently while still achieving the globally optimal outcome macglashan2022value ; russell2003q . However, in the setting studied in this work, this condition does not hold due to a coupling term that depends on how many agents select the crowded action. Specifically, the global reward takes the form
@@
$\hat{r}(s,a;\theta)$ $=\sum_{i}u_{i}(s_{i})+\mathbb{I}_{a=a_{\text{crowd}}}(a_{i})\cdot d\left(1+\sum_{j\neq i}\theta_{j}\right)$ (6)
@@
The following result gives an indication of the error that one could have when considering the approximated definition of the reward in (6). In particular, we define $J(\pi,\beta)$ the discounted reward associated to a policy $\pi$ when considering the reward $r$ and $\hat{J}(\pi,\beta)$ the one obtained when considering the approximated reward $\hat{r}$ defined in (6).
@@
$\mid J(\pi,\beta)-\hat{J}(\pi,\beta)\mid\leq\frac{1}{1-\gamma}\sum_{i}\theta_{i}\left(\frac{\theta_{-i}}{N_{agents}-1}d(N_{agents})-d(1+\theta_{-i})\right)$
@@
$J(\pi_{\text{composed}},\beta)=\hat{J}(\pi_{\text{composed}},\beta)=\sum_{i}\hat{J}_{i}(\pi_{i}^{\star}(\theta^{*}),\beta)$
4.2 Constrained MDP for Each Agent
As introduced in eq. 4, we approximate the global objective by decomposing it into a collection of independent Constrained Markov Decision Processes (CMDPs), one per agent. Constrained Markov Decision Process (CMDP) altman2021constrained extends the standard MDP by incorporating constraints on long-term cost signals. In this section, we describe the structure of the CMDP associated with each agent $i$.
@@ 4.2 Constrained MDP for Each Agent
The agent receives a reward according to the approximated formulation described in section 4.1, where the influence of other agents is captured through the fixed (in the slow and intermediate timescale) parameter $\theta_{-i}$:
$\hat{r}_{i}(s,a;\theta_{-i})=u_{i}(s_{i})+\mathbb{I}_{a_{i}=a_{\text{crowd}}}\cdot d\left(1+\theta_{-i}\right).$
@@
4.3 Learning the Optimal Local Constrained Policy
@@
In particular, we follow the structure of RCPO tessler2018reward for constrained optimization, employing either PPO schulman2017proximal , DQN mnih2013playing or Q-learning watkins1992q , depending on the scale of the system, as the algorithm to handle the fastest timescale
@@
$\hat{r}_{i}^{\text{total}}(s_{i},a_{i};\theta_{-i})=\hat{r}_{i}(s_{i},a_{i};\theta_{-i})+\lambda_{i}c_{i}(s_{i},a_{i}),$ (10)
where $\hat{r_{i}}$ and $c_{i}$ have been defined in section 4.2.
@@
4.4 Coordination via optimization of the constraints
@@
In particular, we want to show that we can converge to the local optimal value of the constraint vector $\theta\in\mathbb{R^{N}}$. The safe RL algorithms described in section 4.3 will then be able to find the optimal policy for that value.
We consider the long term Lagrangian reward with regard to $\theta$, assuming that we can solve the corresponding constrained problem and find the optimal values of $\pi$ and $\lambda$. This corresponds to the quantity we want to minimize for each agent. With abuse of notation, we define the quantity
$\hat{J}_{i}(\theta)$ $=\hat{J}(\pi_{i}^{\star}(\theta))$
$=\mathcal{L}_{i}(\pi_{i}^{\star}(\theta),\lambda_{i}^{\star}(\theta);\theta)$
$=\mathbb{E}_{a\sim\pi^{\star}_{i}(\theta)}\bigg[\sum_{t=0}^{\infty}\gamma^{t}\sum_{t=0}^{T}\hat{r}_{i}(s_{t},a_{t};\theta_{-i})+\lambda_{i}^{\star}\cdot(c_{i}(s_{t},a_{t})-\theta_{i})\bigg]$ (12)
and the function we want to minimize is
$\hat{J}(\theta)=\sum_{i}\hat{J}_{i}(\theta)$ (13)
@@
$\hat{J}_{i}(\theta)$ is differentiable in $\theta$ almost everywhere.
@@
The proof of this result is provided in section A.3, where we show that a variation of the Envelope Theorem milgrom2002envelope can be applied within the context of our setting. ∎
In order to discuss the optimization of the virtual constraint we want to use the stochastic gradient ascent methods of the Kiefer-Wolfowitz family kushner1978stochastic . The iteration scheme is
@@
$\left(\hat{g}_{n}\right)_{i}=\frac{\hat{J}\left(\theta+c_{n}\left(\Delta_{n}\right)_{i}\right)-\hat{J}\left(\theta\right)}{c_{n}(\Delta_{n})_{i}}$ (15)
@@
Then the algorithm converges to a local optimal value of $\min_{\theta}\sum_{i}\hat{J}_{i}(\theta)$.
@@
Lemma 2 guarantees the differentiability of the objective function. The assumptions on the sequences $\alpha_{n}$ and $c_{n}$ allow us to use Proposition 1 in l1994stochastic for a subset of $\Theta$ which contains the initial vector $\theta_{0}$. This proves the convergence of the algorithm to a locally optimal solution. ∎
See Proposition 1 in l1994stochastic for the statement and the necessary conditions for the convergence to the global optimal point.
4.5 Numerical optimization of the constraints via evaluations of the local policies
@@
$\frac{\partial}{\partial\theta_{i}}\hat{J}(\theta_{i},\theta_{-i})=\frac{\partial\hat{J}_{i}(\theta_{i},\theta_{-i})}{\partial\theta_{i}}+\sum_{j\neq i}\frac{\partial\ \hat{J}_{j}(\theta_{i},\theta_{-i})}{\partial\theta_{-j}}$ (16)
@@
Equation 16 illustrates how the problem structure, together with the form of the constrained value function, can be leveraged to compute the gradient of the approximate objective function efficiently. Specifically, this result allows each component of the gradient at a given point $\theta$ to be estimated using only three stochastic evaluations per agent, rather than $N+1$, as we would only need to evaluate $\hat{J}_{i}(\theta_{i},\theta_{-i}),\hat{J}_{i}(\theta_{i}+\epsilon,\theta_{-i})$ and $\hat{J}_{i}(\theta_{i},\theta_{-i}+\epsilon)$.
While Equation 16 enables an efficient numerical estimation of the gradient, the following result provides a more precise analytical characterization of its components and clarifies their qualitative behavior.
Proposition 3.
The following is verified:
  1. 1.
    $\frac{\partial\hat{J}_{i}(\theta)}{\partial\theta_{i}}=-\frac{\lambda^{\star}}{M}\leq 0$ a.e., where $\lambda^{\star}$ is the optimal value of the Lagrange multiplier in $\theta_{i}$ and $M$ is a positive constant,
  2. 2.
    $\frac{\partial\hat{J}_{i}(\theta)}{\partial\theta_{j}}=\theta_{i}\frac{\partial}{\partial\theta_{j}}d(1+\theta_{-i})\geq 0,\ \forall j\neq i$ a.e., where $\pi^{\star}_{i}$ is the optimal policy for agent $i$.
Proof.
Proposition 3 formalizes the intuition about the sign of the gradient components and provides the exact expression for the local derivatives. In particular, relaxing the constraint for one device decreases its own reward (i.e., benefits it from a selfish perspective) while penalizing the others, as a consequence of the shared congestion effect.
It is worth noting, however, that this expression still relies on the exact values of the optimal multipliers $\lambda_{i}^{\star}$, which are not directly available during learning. The proposed algorithm converges to these values asymptotically, but in practice the gradients are approximated. For this reason, in the numerical experiments we currently rely on finite-difference estimates.
The additional simplification obtained by directly substituting $\lambda_{i}^{\star}$ in place of the local finite-difference evaluations could further reduce computation times. This refinement has not yet been tested in the stochastic approximation setting, though we have verified its correctness in small systems where $\lambda_{i}^{\star}$ was computed exactly via linear programming (section 5.5). Evaluating the performance of this direct substitution in the approximate setting is left for future work.
@@
5 Numerical experiments
5.1 Scalability properties
5.2 Offloading action frequency
5.3 Effect of Initial Constraint
5.4 Non linear penalty
5.5 Evaluation gradient objective function
Additional experimental results are provided in the appendix, including the complete data underlying fig. 9, the evolution of offloading behavior in larger systems, comparisons with alternative baselines, and numerical evaluations of the theoretical results from lemma 1 using linear programming. Together, these supplementary experiments reinforce the conclusions drawn from the main results.
6 Conclusions
Appendix A Theoretical proofs
A.1 Proof of Lemma 1
@@
$J(\pi,\beta)-\hat{J}(\pi,\beta)$ $=\mathbb{E}_{a_{t}\sim\pi}\left[\sum_{t}\gamma^{t}\sum_{i}r_{i}(s_{t},a_{t})\mid s_{0}\sim\beta\right]-\mathbb{E}_{a_{t}\sim\pi}\left[\sum_{t}\gamma^{t}\sum_{i}\hat{r}_{i}(s_{t},a_{t})\mid s_{0}\sim\beta\right]$
@@
$J(\pi,\beta)-\hat{J}(\pi,\beta)\leq\frac{1}{1-\gamma}\sum_{i}\theta_{i}\left(\frac{\theta_{-i}}{N_{agents}-1}d(N_{agents})-d(1+\theta_{-i})\right)$
@@
$\left|J(\pi,\beta)-\hat{J}(\pi,\beta)\right|\leq\frac{1}{1-\gamma}\sum_{i}\theta_{i}\left(\frac{\theta_{-i}}{N_{agents}-1}d(N_{agents})-d(1+\theta_{-i})\right)$ (18)
A.2 Theoretical background for the proof of lemma 2